Two slits spaced 0.130 mm apart are placed 90.0 cm from a screen and illuminated by coherent light with wave length 550 nm. the intensity at the center of the central maximum (Two slits spaced 0.130 mm apart are placed 90.0 cm degrees) is 4.00 X 10-6 W/m2 A) what is the distance on the screen from the center of the central maximum to the first minimum? B) what is the intenisty at a point midway between the center of the central maximum and the first minimum?
Answer #1

Two slits spaced (2d) = 0.130 mm =0.130*10-3m

The distance between the slits and screen is (D) = 90.0 cm =0.9m

The illuminated by coherent light with wave length (lamda) = 550 nm =550*10-9m.

The intensity at the center of the central maximum (\theta =0 degrees) is (I) =4.00 X 10-6 W/m2

For a first dark fringe i.e fo the first minima

                                    xn =(D/2d)(2n+1)lamda/2      for firstminima n=1

now the distance on the screen from the center of the central maximum to the first minimum

                                   xn =(D/2d)*(2n+1)lamda/2 =(0.9/0.130*10-3m)(3)*550*10-9m/2 =571.15*10-6 =5.711mm

At the mid point means

                                    xn =2.855mm